2r2-3ry equation for the tangent plane to the graph of f(x, y) (2, 1). 5. Find an at (r, y)
Answer #1

3 4(x) -3 2.1) 8-3 92x1 5,-1 8 (2, 1) 2 6 = 7 ea tagent plane: = (1-19.F18) Cn-2).S t 2 + 2+ 574-10 8T+ Rer-+ 5n-18+10 Z= 5